### v2

# When Models Outthink Their Safety: Mitigating Self-Jailbreak in Large Reasoning Models with Chain-of-Guardrails

Abstract
Large Reasoning Models (LRMs) demonstrate remarkable capabilities on complex reasoning tasks but remain vulnerable to severe safety risks, including harmful content generation and jailbreak attacks. Existing mitigation strategies rely on injecting heuristic safety signals during training, which often suppress reasoning ability and fail to resolve the safety-reasoning trade-off. To systematically investigate this issue, we analyze the reasoning trajectories of diverse LRMs and uncover a phenomenon we term Self-Jailbreak, where models override their own risk assessments and justify responding to unsafe prompts. This finding reveals that LRMs inherently possess the ability to reject unsafe queries, but this ability is compromised, resulting in harmful outputs. Building on these insights, we propose the Chain-of-Guardrail (CoG), a training framework that recomposes or backtracks unsafe reasoning steps, steering the model back onto safe trajectories while preserving valid reasoning chains. Extensive experiments across multiple reasoning and safety benchmarks demonstrate that CoG substantially improves the safety of current LRMs while preserving comparable reasoning ability, significantly outperforming prior methods that suffer from severe safety–reasoning trade-offs.
1 Introduction
Large Reasoning Models (LRMs), such as OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), and Qwen3 (Yang et al., 2025), have demonstrated remarkable capabilities in solving complex tasks and are increasingly being adopted across a wide range of domains. Despite these advances, researchers have identified that current LRMs remain hindered by significant safety vulnerabilities, such as the generation of harmful content as well as the susceptibility to jailbreak attacks (Zhou et al., 2025a; Zhang et al., 2025; Wang et al., 2025a). These issues highlight a critical obstacle to the reliable and responsible use of LRMs (Weidinger et al., 2021). Therefore, addressing these challenges is critical for ensuring that LRMs operate safely and remain aligned with human values.
While recent efforts (Wang et al., 2025b; Jeung et al., 2025; Zhou et al., 2025a; Jiang et al., 2025) have attempted to mitigate these risks, they largely exhibit a pronounced safety-reasoning trade-off. Specifically, current studies have primarily focused on curating safety-oriented datasets based on heuristic rules and fine-tuning LRMs to produce refusal responses. Unfortunately, while such approaches can improve safety to some extent, the pattern-based behavior often conflicts with the flexible and multi-step deliberation required for complex reasoning, thereby substantially degrading reasoning ability. Therefore, it is critical to identify the underlying causes of the unsafe behavior in LRMs, and develop more principled strategies that can enhance safety without sacrificing reasoning performance.
To this end, we conduct a systematic analysis of LRMs’ reasoning trajectories and identify a critical failure phenomenon, which we term Self-Jailbreak. Specifically, as demonstrated in Figure 1, we decompose LRM’s reasoning thinking trajectory into three stages: Risk Awareness, Risk Analysis, and Response Strategy. We observe that LRMs often detect harmful intent during the Risk Awareness stage, but subsequently override this recognition during the Risk Analysis stage, and ultimately persuade themselves to respond to unsafe queries. This reveals that LRMs possess an inherent ability to reject unsafe queries, yet it is undermined by Self-Jailbreak, resulting in unsafe generation. Furthermore, we categorize Self-Jailbreak into four distinct types—Benign Reframing, Logical Fallacy, Warning, and Risk Misrecognition—and find that Warning accounts for most unsafe outputs.
Existing approaches that solely inject heuristic safety signals during training fail to resolve the safety–reasoning trade-off. While they suppress overtly unsafe outputs, such interventions do not address unsafe reasoning steps embedded within reasoning trajectories. Building on these insights, we propose Chain-of-Guardrail (CoG), a training framework designed to mitigate Self-Jailbreak while preserving reasoning quality. Specifically, CoG leverages precise classification of Self-Jailbreak to perform targeted safety interventions, mitigating unsafe behavior while preserving the model’s reasoning abilities and reasoning patterns. To achieve this, CoG employs two strategies: 1) Safety Recomposition, which systematically reorganizes unsafe reasoning into entirely new, safety-assured chains, and 2) Safety Backtrack, which preserves the model’s original reasoning path while incorporating self-check and backtracking mechanisms to correct risky steps before they lead to unsafe conclusions. The resulting safety-oriented reasoning chains are then used to fine-tune LRMs, aligning them toward safe responses while maintaining the reasoning capability.
Figure 1: Illustration of “Self-Jailbreak.” The Instruct model (left) directly refuses harmful queries, whereas the LRM (right) identifies risk during Risk Awareness but later engages in Self-Jailbreak (red text) during Risk Analysis, ultimately overriding its initial safety judgment and producing harmful content in the Response Strategy stage.
Extensive experiments across multiple safety and reasoning benchmark datasets demonstrate that COG excels at balancing the safety and reasoning performance of LRMs. On Qwen3-32B, CoG improves the reasoning ability by 13.5% over SafeKey while achieving state-of-the-art performance on GPQA and AIME (Rein et al., 2024; Mathematical Association of America (2024), MAA), maintaining comparable reasoning performance. Further analysis of reasoning trajectories reveals that, unlike other approaches, CoG preserves the model’s intrinsic reasoning patterns, confirming that its safety gains do not come at the expense of the model’s original deliberation trajectory. In summary, our contributions include: In summary, our contributions include:
  • To the best of our knowledge, we are the first to identify and characterize the Self-Jailbreak phenomenon, revealing it as a key factor behind the poor safety performance of LRMs.
  • We propose a systematic analytical framework to categorize Self-Jailbreak instances and quantify their impact on model behavior, finding that Benign Framing and Warning constitute the majority of the Self-Jailbreak phenomenon.
  • We introduce the Chain-of-Guardrail (CoG), a training framework that mitigates Self-Jailbreak and substantially improves model safety while preserving reasoning capabilities. Experiment results show that CoG improves the safety-reasoning trade-off by 3.39%, reaching the state-of-the-art (SOTA) across multiple LRMs.
Figure 2: Processing of harmful queries by LRM, illustrating the workflow from receiving prompts, through risk detection during reasoning, to the four possible outcomes: aware of risk and responds, aware of risk and refuses, unaware of risk and responds, unaware of risk and refuses.
2 Motivation: Preliminary Experiments On Self-Jailbreak
This section examines the thinking trajectory of LRMs, drawing on pilot experimental results (reported in Appendix B.1) which indicate that most LRMs become unsafe when the thinking mode is enabled. Subsequently, we find that even if LRMs are capable of recognizing latent dangerous intentions, they still respond to the harmful query. Such results shed light on our hypothesis: Self-Jailbreak. We further validate the existence of Self-Jailbreak with experiments, which represent two significant findings: (1) Most LRMs encounter Self-Jailbreak during the reasoning trajectory. (2) Among the different categories of Self-Jailbreak, Warning is the primary source.
2.1 Hypothesis: Self-Jailbreak via Thinking–Answer Misalignment
We analyzed the models’ chains of thought, discovering that their thinking trajectory can be decomposed into three stages: risk awareness, risk analysis, and response strategy, which enables a deeper examination of how thinking affects the safety of LRMs.
Building on the previous experiment, we employ LLM-as-Judge (Team, 2024; Gu et al., 2024) to assess whether the LRM demonstrates awareness of potential risks during the risk awareness stage, and use Llama-Guard (Llama Team, 2024) to evaluate whether the final response contains harmful content. This setup enables us to examine potential misalignments between the model’s reasoning trajectory and its generated answers. As illustrated in Figure 2, our analysis revealed a striking contradiction:
LRMs are able to identify most of the harmful queries but still generate unsafe answers. 94.2% of harmful queries, as well as users’ potentially malicious intents, are recognized; however, 52.30% of their generated outputs still contain hazardous content. In contrast, only 39.90% of the cases reflect awareness of risk followed by a safe response. This inconsistency between “thinking” and “response” raises a critical concern: during subsequent reasoning, LRMs may tend to persuade themselves to respond to such queries. Based on this observation, we propose the following hypothesis:
Self-Jailbreak During the thinking process, LRMs often exhibit a phenomenon we term Self-Jailbreak, where models persuade themselves to produce responses to potentially harmful requests in an attempt to assist the user, even after recognizing the risky nature of the queries and the user’s underlying malicious intent.
Table 1: Self-Jailbreak category distribution of various LRMs (%)
Model Response Safety Self-Jailbreak Categories
Safe Unsafe Benign Reframing Warning Logical Fallacies Harm Identification
DS-R1-0528 4.07 95.93 42.29 52.12 0.51 1.02
DS-Llama-70B 0.50 99.50 47.89 45.67 0.91 5.04
Qwen3-8B 0.21 99.79 38.60 50.69 2.41 8.10
Qwen3-14B 0.15 99.85 37.30 53.10 1.73 7.32
Qwen3-32B 0.36 99.64 36.32 52.51 2.20 8.61
2.2 Categorization and Existence of Self-Jailbreak
Building on this hypothesis and inspiration from backtracking during mathematical reasoning  (Chen et al., 2025), we designed a more rigorous classification framework, which is guided by the principle that the model’s reasoning trajectory should not generate content that directly or indirectly facilitates harmful user behaviors. We categorize these behaviors into four types:
Benign Reframing:
The model recognizes the risk but attempts to reinterpret the user’s intent benignly, offering help from such interpretation.
Warning:
The model assumes that appending a warning to harmful content is sufficient to prevent its misuse, and proceeds to generate its answer.
Logical Fallacies:
Given a query containing logical contradictions, the model detects potential harm but falls into logical traps, leading to logical fallacies.
Harm Identification:
The model fails to recognize the harm in the query, resulting in answering.
To investigate the prevalence of Self-Jailbreak in LRMs, we categorize and analyze the thinking trajectory of several contemporary LRMs on 2k queries in WildJailbreak (Jiang et al., 2024a). As shown in Table 1, we observe that:
(1) Self-Jailbreak is a prevalent phenomenon observed in various LRMs. Experiment results indicate that Self-Jailbreak occurs frequently across LRMs, including peripheral models like DeepSeek-R1-0528 (DeepSeek-AI, 2025), highlighting the importance and relevance of our study.
(2) Warning is the most common source of Self-Jailbreak. As illustrated in Table 1, Warning is the most categorized Self-Jailbreak type across various LRMs. This tendency likely stems from the fact that LRMs are largely trained on math and code data, which predisposes them to respond in ways that meet user expectations. Consequently, even when confronted with harmful queries, LRMs are likely to be helpful, inadvertently bypassing their own safety mechanisms by providing content alongside a safety disclaimer. Such behavior ultimately gives rise to Self-Jailbreak.
3 Chain-of-Guardrail: A Training Framework to Mitigate Self-Jailbreak
To systematically address the risks of Self-Jailbreak in LRMs, we propose a unified safe training framework: Chain-of-Guardrail (COG). As illustrated in Figure 3, the framework is designed to create an interpretable and structured pipeline that automatically removes or rejects unsafe content during the reasoning trajectory $c$, producing a safe and logically consistent response $y_{\text{safe}}$. To achieve this, COG employs two complementary methods: Safety Recomposition (SafR) and Safety Backtrack (SafB). SafR focuses on proactively restructuring and correcting the reasoning chain to eliminate potential safety risks, while SafB focuses on the reasoning trajectory when unsafe content is detected, ensuring that the model corrects errors and generates a safe response.
[…]
3.1 Optimization Goal of LRMs in COG
Specifically, we formulate the optimization objective as follows:
$f:(\pi_{0},x)\rightarrow y_{\text{safe}},\quad\text{s.t.}\quad\text{Preserve}(\text{Reasoning}_{\pi_{0}})$ (1)
where $\pi_{0}$ is the original model, $y_{\text{safe}}$ is the safe response, and $\text{Preserve}(\cdot)$ ensures maximal retention of the original reasoning.
To achieve this, we explicitly model the Chain-of-Thought (CoT) as a structured latent variable:
$c=\pi_{0}(x)=[d(x),a(x),p(x)],$ (2)
where $d(x)$, $a(x)$, and $p(x)$ represent the risk awareness, risk analysis, and response strategy.
Depending on the adopted process, the safety-oriented reasoning chain is constructed as follows:
Chain of Safety Recomposition (SafR):
$c_{\text{safe}}=\text{Merge}(d(x),\hat{a}(x),\hat{p}(x)),$ (3)
where $\hat{a}(x)$ and $\hat{p}(x)$ are the risk analysis and response strategy recomposed by the Safety Recomposition procedure.
Chain of Safety Backtrack (SafB):
$c^{\prime}_{\text{safe}}=[c,s(x)],$ (4)
where $s(x)$ denotes the model-generated self-check module appended to the original reasoning chain in the Safety Backtrack procedure.
The final safe response $y_{\text{safe}}$ is produced based on the safety-oriented reasoning chain derived from either SafR or SafB.
3.2 Selective Loss Masking Strategy
To avoid undesired shifts in the model’s predicted distribution due to safety interventions, we employ a selective loss masking strategy. Let the output sequence be $o=[c_{\text{safe}},y_{\text{safe}}]$ for SafR, and $o=[c^{\prime}_{\text{safe}},y_{\text{safe}}]$ for SafB, where $c_{\text{safe}}$ and $c^{\prime}_{\text{safe}}$ are defined as in Eq. 3 and Eq. 4, respectively.
We propose the utilization of a binary mask vector $m\in\{0,1\}^{|o|}$, to supervise distinct segments of the output sequence $o$ selectively. The primary training goal can be summarized as follows:
$\mathcal{L}_{SafR/SafB}=-\sum_{i=1}^{|o|}m_{i}\cdot\log P_{\theta}(o_{i}|o_{<i},x),$ (5)
This selective loss masking strategy allows for flexibility depending on the applied process:
Safety Recomposition: $m$ covers all tokens, optimizing both the revised reasoning chain and the safe response, allowing the model to learn the complete safe reasoning trajectory.
Safety Backtrack: $m$ covers only the appended self-check and final response, while the original chain is masked as context, preserving the model’s reasoning distribution and enhancing self-reflection and safety.
3.3 Workflow of COG
Both SafR and SafB aim to enhance safety by mitigating certain aspects of Self-Jailbreak, following a three-phase structure. While their core workflows overlap, they differ in how the reasoning chain is modified and integrated.
Phase 1: Input & Analysis (shared)
Initial Inference. The base model $\pi_{0}$ infers on the harmful prompt $x$, producing the initial reasoning chain $c=[a(x),p(x)]$ and response $y$.
Information Extraction. A structured prompt (see Appendix A.2) together with Qwen2.5-72B-Instruct (Team, 2024) extracts the risk analysis and response strategy.
Failure Classification. Given $x$, $a(x)$, and $p(x)$, Qwen2.5-72B-Instruct determines whether the reasoning violates safety principles and assigns a Self-Jailbreak type.
Phase 2: Data Construction (divergent)
SafR – Reasoning Chain Recomposition and Merging. For each identified failure type, a tailored prompt guides $\pi_{0}$ to perform reasoning chain recomposition, adjusting the risk analysis and response strategy with safety-oriented considerations to generate corrected sub-chains. Subsequently, these recomposed analyses $a(x)$ and predictions $p(x)$ are merged into a coherent, Safety-oriented Chain-of-Thought (S-COT).
SafB – Safety Backtrack Augmentation. Rather than recomposing, $\pi_{0}$ performs a reflective process on the original reasoning chain to generate a supplementary safety check $s(x)$. Carefully engineered transitional expressions ensure a natural integration between the reasoning chain and the check, resulting in a coherent S-COT.
Phase 3: Inference & Alignment (shared, with minor variation)
Safe Response Generation. Both SafR and SafB use the S-COT as the reasoning trajectory to guide $\pi_{0}$ in generating a safe response $y_{\text{safe}}$.
Alignment Training. In both methods, $\pi_{0}$ is further trained on the safe response data, yielding the aligned and safety-enhanced model $\pi_{1}$ with strong reasoning capabilities.
4 Experiment
4.1 Experiment Setting
Baselines We include four representative methods: STAR-1  (Wang et al., 2025b) and SafeChain  (Jiang et al., 2025) improve safety by performing SFT on a curated set of safe data. SAFEPATH  (Jeung et al., 2025) guides model behavior by training it to generate a fixed prefix, and SafeKey  (Zhou et al., 2025b) enforces stricter safety alignment by introducing two additional safety classifiers. Each method was reproduced using the officially released training configurations and datasets, with details provided in Appendix A.1.
Models and Configuration To validate the efficacy of our methods, we utilize the Qwen3 series  (Yang et al., 2025). Our Safety Recomposition and Safety Backtrack methods are implemented via supervised fine-tuning (SFT) on the Qwen3 base models. As illustrated in Figure 3, the training and experimental procedures comprise three distinct phases. The details are shown in the Appendix A.2.2.
Safety Benchmarks We evaluate model safety using four benchmarks. SORRY-Bench  (Xie et al., 2024) and StrongREJECT  (Souly et al., 2024) measure the model’s ability to refuse harmful prompts, while WildJailBreak  (Jiang et al., 2024b) and JailBreakBench  (Chao et al., 2024) assess its robustness against adversarial jailbreak attacks.
[Uncaptioned image] Figure 4: Safety vs. reasoning trade-off for the 32B model.
Reasoning Benchmarks To assess general reasoning ability, we use AIME2024  (Mathematical Association of America (2024), MAA) for mathematical reasoning and GPQA-Diamond  (Rein et al., 2024) for complex knowledge-based reasoning. To reduce randomness in evaluation, we perform multiple rollout rounds: GPQA is averaged over two rounds, and AIME over eight rounds. Additional details of the reasoning and safety benchmarks can be found in Appendix A.3.
Training Dataset We extracted 15,000 high-quality harmful prompts from 7 public collections—Alert  (Tedeschi et al., 2024), ToxicDPOqa, Harmful-Dataset, Aya_RedTeaming  (Ahmadian et al., 2024), Do-Not-Answer  (Wang et al., 2023), AttaQ  (Kour et al., 2023), and Toxic-Chat  (Lin et al., 2023). We then designated this curated set as our seed dataset and fed it into our pipeline to generate corresponding high-quality safety response sets.
4.2 Main Result
Table 2: Performance comparison of different methods under the main experimental setting. Here, lower values in harmful/jailbreak benchmarks and higher values in reasoning benchmarks represent better performance.
Method Harmful Jailbreak Reasoning
[…]
Vanilla 45.45 13.62 38.80 81.71 57.33 77.50
STAR-1 18.86 0.74 80.00 62.2 57.33 71.25
SAFEPATH 36.14 10.03 22.80 42.68 55.56 67.92
SafeChain 49.55 16.99 36.80 70.73 52.28 66.25
SafeKey 3.18 0.32 8.53 13.2 41.90 70.80
Safety Backtrack(Ours) 17.05 2.05 8.00 26.83 54.30 77.50
Safety Recomposition(Ours) 13.41 2.86 9.20 28.05 56.82 76.25
[…]
Vanilla 55.45 12.44 34.00 68.29 63.14 77.92
STAR-1 17.95 0.72 86.80 76.83 56.32 76.25
SAFEPATH 34.09 8.49 16.00 50.00 57.78 70.42
SafeChain 49.32 16.87 35.60 67.07 57.58 71.25
SafeKey 4.77 0.32 4.88 10.40 49.00 76.70
Safety Backtrack(Ours) 11.82 1.30 6.40 18.05 62.12 77.92
Safety Recomposition(Ours) 12.50 4.97 2.80 21.95 60.36 78.75
[…]
Vanilla 46.59 12.25 35.20 20.43 65.66 81.67
STAR-1 18.41 0.74 16.80 35.37 54.55 72.92
SAFEPATH 40.00 6.57 22.80 53.66 62.38 70.25
SafeChain 47.95 16.39 28.40 70.73 54.30 71.70
SafeKey 3.41 0.32 7.32 10.40 54.30 71.70
Safety Backtrack(Ours) 14.55 2.77 8.80 23.17 61.62 77.08
Safety Recomposition(Ours) 7.05 1.49 3.20 17.07 62.38 82.08
Our results, summarized in Table 4.2, reveal several key findings:
Baseline methods enhance safety performance at the substantial cost of reasoning capability. As reported in Table 4.2, several baselines either underperform the vanilla model on safety (e.g., SafeChain, STAR-1) or secure safety gains only at the cost of substantial reasoning degradation (e.g., SafeKey).
COG achieves the best trade-off between safety and reasoning across model scales. Our proposed COG framework, including Safety Backtrack and Safety Recomposition, consistently achieves the strongest safety improvements while largely preserving reasoning ability. This trade-off is clearly visualized in Figure 4, where our methods are positioned in the optimal upper-right quadrant, demonstrating both high safety and high reasoning performance simultaneously. Specifically, across model scales, SafB and SafR attain the lowest or near-lowest harmful/jailbreak scores while maintaining competitive reasoning performance (GPQA-Diamond and AIME2024), highlighting the robustness and scalability of COG in resolving the safety–reasoning trade-off.
5 Analysis
5.1 PCA Analysis of Safety Improvements with Minimal Reasoning Impact
To quantitatively evaluate the trade-offs between safety and reasoning, we apply Principal Component Analysis (PCA) to the hidden states of Qwen3-32B We analyze the geometric structure of the representation space under different training approaches, focusing on safety-critical prompts (Harmful, RedTeaming) and reasoning prompts (Reasoning). A linear decision boundary is trained on the two-dimensional PCA projection to quantify the separation between safety and reasoning  [^1] Results for the 8B and 14B models are provided in Appendix B.2., which is illustrated in Figure 5, where we can conclude that:
Our approach establishes a more separable safety manifold while retaining the stability of reasoning representations, indicating effective safety gains with negligible reasoning degradation. Specifically, SafR achieves the largest safety improvement ($\Delta$_Harmful = +4.985) with minimal reasoning loss ($\Delta$_Reasoning = –0.676), whereas SafB provides a more balanced gain ($\Delta$_Harmful = +3.961, $\Delta$_Reasoning = –1.771). In contrast, SafeChain offers only marginal safety benefits ($\Delta$_Harmful = +0.378) while causing severe reasoning degradation ($\Delta$_Reasoning = –5.260). These results indicate that our COG framework substantially enhances safety while largely retaining reasoning capabilities, effectively improving safety alignment without disrupting core reasoning structures.
Figure 5: PCA of the Qwen3-32B representation space. Gray ellipses indicate base model distributions, while colored regions denote post-training clusters (red: safety, blue: reasoning). Centroid shifts ($\Delta$) quantify displacement relative to the linear decision boundary (dashed), with $\Delta>0$ indicating increased margin and $\Delta<0$ indicating decreased margin.
5.2 Understanding Reasoning Preservation via Trajectory and Token Analysis
To study how training strategies influence the reasoning structure of language models, we analyze their cognitive trajectories using DeepSeek-v3.1 (DeepSeek-AI, 2024) with established prompts (Zeng et al., 2025; Gandhi et al., 2025), and measure how these patterns differ from a base model. We track the frequency of key reasoning behaviors—Backtracking, Enumeration, Subgoal Setting, and Verification—with scores reflecting the average occurrences per problem on the AIME benchmark, which is presented in Table 3, which indicates that:
Models trained with our methods best maintain the cognitive structure of the base model, and therefore have only a minimal impact on the model’s reasoning capabilities. Specifically, SafR exhibits slightly higher behavior frequencies (+0.3%) and SafB slightly lower (–0.3%), indicating that both methods achieve high alignment with the base model’s reasoning patterns. In contrast, alternative methods such as SafeChain, SafePath, and star-1 exhibit larger deviations (see Table 3). Furthermore, the token-level analysis in Appendix B.3 confirms that SafR and SafB maintain token usage close to the Base model while improving safety, indicating that they enhance safety without significantly altering intrinsic reasoning behavior.
Table 3: Comparison of reasoning patterns (Qwen3-32B) across different training strategies.“Overall Avg.” denotes the average frequency across all reasoning patterns, reflecting the overall reasoning style shift under different strategies.
Pattern Base SafeChain SafePath star-1 SafB SafR
Backtracking 1.33 1.10 1.20 1.30 1.27 1.30
Enumeration 0.93 0.87 0.97 0.83 1.00 1.03
Subgoal Setting 1.60 1.63 1.30 1.40 1.47 1.57
Verification 2.50 2.47 2.23 2.10 2.50 2.57
[…]
Table 4: Comparison of model performance with and without masking the default reasoning trajectory. “SafB (full mask)” corresponds to the Safety Backtrack (full mask) setting, while “SafB (part. mask)” corresponds to the Safety Backtrack (partial mask) setting.
[…]
Qwen3-8B 45.45 13.62 38.80 81.71 57.58 73.30
SafB (full mask) 17.05 2.05 8.00 26.83 54.30 77.50
SafB (part. mask) 23.64 5.37 22.40 59.76 56.87 75.00
5.3 Impact of Masking Strategy in Safety Backtrack
We evaluate the masking strategy in Safety Backtrack (SafB) by comparing two training setups: (1) Full-mask training, where the loss is applied to the reasoning trace, self-check, and final answer; and (2) Partial-mask training (our main method), where the loss is applied only to the self-check and final answer, with the reasoning trace masked.
Table 4 shows that both SafB variants improve safety, with partial masking yielding larger gains on the Harmful and RedTeaming benchmarks. For reasoning, both variants enhance AIME, whereas GPQA shows a slight degradation, more pronounced under full masked SafB, which can be attributed to fewer rollout samples (two for GPQA vs. eight for AIME).
6 Related Work
Safety–Reasoning Trade‐Off
Reasoning capability and safety appear to be inherently conflicting objectives in LRMs. Zhou et al.  (Zhou et al., 2025a) show that stronger LRMs tend to produce more unsafe content, while studies by Huang et al.  (Huang et al., 2025) and Li et al.  (Li et al., 2025) highlight a fundamental trade‐off: improving one objective inevitably undermines the other, underscoring the need to balance safety and reasoning performance.
Safety for Reasoning
Instead of relying on external filters, safety-for-reasoning methods integrate controls directly into inference through training. Current approaches mostly build on supervised fine-tuning (SFT). STAR-1  (Wang et al., 2025b), SafeChain  (Jiang et al., 2025), SAFEPATH  (Jeung et al., 2025), and SafeKey  (Zhou et al., 2025b) apply curated data and tailored SFT to enhance safety while retaining reasoning ability. These methods better preserve coherence but still deliver only modest safety gains or impose notable performance costs, leaving room for more efficient optimization.
7 Conclusion
In this paper, we conducted a systematic investigation into the safety mechanisms of LRMs. Our analysis reveals the existence of Self-Jailbreak, where LRMs persuade themselves into answering the harmful queries. This phenomenon sheds new light on the underlying mechanisms of the reasoning process in LRMs. Building on this discovery, we developed CoG, a training framework that suppresses Self-Jailbreak by detecting and mitigating unsafe reasoning patterns through Safety Backtrack and Safety Recomposition. Experimental results show that CoG substantially enhances safety with minimal degradation in reasoning capability, achieving state-of-the-art performance in the safety–reasoning trade-off. Overall, this study provides a principled framework for aligning LRMs toward safe and reliable reasoning.
Appendix A Detailed Experimental Setup
A.1 Baseline Configuration
A.1.1 Description of Baseline Methods
STAR-1
STAR-1 categorizes 41,000 safety data points from multiple sources into eight predefined categories and generates a response with CoT using DeepSeek-R1, guided by the safety policies associated with each category. Then, a set of rules is applied to filter out 1,000 data points for the dataset. These data are then used to fine-tune an LRM to conduct safety alignment.
SafeChain
SafeChain selected 50,000 data points from the wildjailbreak dataset and used R1-70B to generate five responses for each instruction. Then, Llama-Guard is used to filter data, keeping the responses that are all safe. A random response is sampled from the five responses as the final response. This created a dataset containing 40,000 instruction-response pairs, available for supervised fine-tuning.
SAFEPATH
SAFEPATH fine-tunes LRMs in a specific way, making them always generate eight fixed tokens: “let’s think about safety first” at the start of inference, guiding the LRMs to consider more about safety during the generation process.
SafeKey
SafeKey enhances safety reasoning by integrating a Dual‑Path Safety Head with Query‑Mask Modeling to amplify latent safety signals from both the raw input (X) and the model’s internal query understanding (U) during generation of the “key sentence”—this effectively triggers a safety-focused “Aha moment.” By masking out X when predicting the key sentence based solely on U, Query‑Mask Modeling strengthens the U→K pathway, while the dual‑path head reinforces these hidden‐state safety cues during fine‑tuning. Together, these two jointly improve robustness against harmful prompts.
A.1.2 Implementation Details of Baselines
Computational Resource
To ensure fair comparison and reproducibility, all experiments—including those reproducing related work—were performed on 8 A-800 with bf16 precision enabled, which allows for faster training while preserving numerical stability. The corresponding training hyperparameters are summarized as follows.
Star-1
We use the official dataset and replicate the experiments following the parameter settings reported in the original paper. The detailed training configurations are presented in Table 5.
Table 5: Detailed Training Hyperparameters for Star-1
Hyperparameter Value Hyperparameter Value
Finetuning Type Full Optimizer AdamW
Adam $\beta_{1},\beta_{2}$ 0.9, 0.95 Learning Rate 1e-5
Epochs 5.0 Batch Size 2
Gradient Accumulation Steps 8 Weight Decay 1e-4
Warmup Ratio 0.05 Cutoff Length 8,192
SafeChain
We trained Qwen3 series models with the original SafeChain dataset with llama-factory. Detailed implementation of SafeChain experiment is described as shown in Table 6:
Table 6: Detailed Training Hyperparameters for SafeChain
Parameter Value Parameter Value
Epochs 2 Batch Size 2
Gradient Accumulation Steps 2
Safekey
We use the official SafeKey codebase, making only model-level modifications to its startup scripts. The detailed implementation of the SafeKey experiment is described as shown in Table 7.
Table 7: Detailed Training Hyperparameters for SafeKey
Parameter Value Parameter Value
Epochs 5 Batch Size 2
Gradient Accumulation Steps 8
SafePath
The detailed implementation of the SafePath experiment is described in Table 8:
Table 8: Detailed Training Hyperparameters for SafePath
Parameter Value Parameter Value
Finetuning Type Full Cutoff Length 8192
Batch Size 2 Gradient Accumulation Steps 2
Learning Rate 1e-5 Max Steps 20
Warmup Ratio 0.05
A.2 Implementation Details of Our Method
A.2.1 Manual Classification Accuracy
Table 9: Classification Accuracy of Three Graduate Students on 50 Cases
Participant Correct Cases Accuracy (%)
Graduate Student A 46 92.0
Graduate Student B 43 86.0
Graduate Student C 44 88.0
Average 44.33 88.67
We invited three graduate students with domain expertise to classify 50 representative cases manually. As shown in Table 9, their classification accuracy ranged from 86.0% to 92.0%, with an average of 88.67%, indicating the reliability of our labeling scheme.
A.2.2 COG Generation Parameters
During the sampling process(Phase 1), to ensure output diversity and prevent model degeneration, we set the temperature to 0.7, top_p to 0.8, and presence_penalty to 1.5 to produce the original responses used as seed data (see Table 10).
During the extraction and classification process(Phase 1), temperature and top_p were set to 0.1 and 0.9, respectively, to ensure that the model outputs its most confident predictions.
During the Safety Recomposition and Safety Backtrack stages (Phase 2), we aimed to maintain consistency between generated content and prompt constraints while preserving diversity; thus, temperature was set to 0.3 and top_p to 0.8.
Finally, for the chain-of-thought based response generation stage(Phase 3), temperature was again set to 0.7, top_p to 0.8, and presence_penalty to 1.5 to maintain diversity.
These carefully chosen parameters balance generation quality and diversity while minimally impacting the model’s reasoning capability.
Table 10: Generation Parameter Settings
Stage Parameters Values
Generation Phase temperature 0.7
top_p 0.8
presence_penalty 1.5
Extraction & Classification temperature 0.1
top_p 0.9
SafR & SafB Phases temperature 0.3
top_p 0.8
Chain-of-Thought Generation temperature 0.7
top_p 0.8
presence_penalty 1.5
Table 11: Training Hyperparameters
Parameter Value Parameter Value
Finetuning Type full Learning Rate 2e-6
Cutoff Length 8192 Epochs 3.0
Batch Size 2 Warmup Ratio 0.1
Gradient Accumulation Steps 4
Table 12: Benchmark Implementation Details
Safety Benchmarks Reasoning Benchmarks
Parameters Values Parameters Values
temperature 0.7 temperature 0.6
top_p 1 top_k 20
max_new_tokens 16384 top_p 0.95
rollout 1 max_seq_length 32768
max_out_len 32000
GPQA rollout 2
AIME2024 rollout 8
COG Training Parameters
Both the Safety Recomposition and Safety Backtrack tasks are trained using LlamaFactory under consistent experimental settings. Our approach is based on a dataset of 14,000 examples, with the full training hyperparameters summarized in Table 11.
A.3 Evaluation Details
A.3.1 Benchmark Description
Sorry-bench
Sorry-bench is a systematic safety-refusal benchmark comprising 440 harmful prompts across 44 fine-grained safety categories. We used the original prompts as the test set to evaluate LLM refusal behaviors.
StrongREJECT
StrongREJECT is a jailbreak robustness benchmark featuring 313 carefully filtered harmful prompts spanning six major misuse categories to assess LLM defenses against jailbreaks.
WildJailbreak
WildJailBreak is an adversarial evaluation split of 2,213 jailbreak prompts drawn from a 262 K-example synthetic safety corpus generated by the WildTeaming framework, designed to rigorously test LLM safety mechanisms. We randomly selected 250 prompts from the evaluation split as the evaluation set.
JailBreakBench
JailBreakBench is a robustness benchmark offering 100 paired harmful-behavior prompts (55 % original, 45 % sourced from AdvBench and TDC/HarmBench). In our experiment, we used harmful prompts augmented with Vicuna-generated PAIR variants for comprehensive jailbreak evaluation.
GPQA-Diamond
GPQA-Diamond is the “Diamond” subset of the GPQA benchmark, comprising the 198 most difficult of 448 graduate-level, domain-expert-written multiple-choice questions in biology, chemistry, and physics.
AIME2024
AIME is the complete set of 30 official integer-answer problems from the 2024 American Invitational Mathematics Examination I & II, directly sourced from the MAA’s public releases.
A.3.2 Evaluation Metrics
For the safety benchmarks, Sorry-Bench, StrongREJECT, and WildJailBreak use attack successful rate (ASR) as the evaluation metric, revealing the times that a model accepts harmful prompts. Following the setting of the original benchmark, we used the rejection rate for JailBreakbench, measuring how often the model successfully rejects harmful prompts. For reasoning benchmarks, we use accuracy as the evaluation metric, measuring the rate at which models give correct answers.
A.3.3 Benchmark Hyperparameters Details
We used a rollout of 2 for GPQA-Diamond and 8 for AIME2024. The evaluations on GPQA-Diamond and AIME2024 were conducted using the OpenCompass framework. The detailed hyperparameter setting is shown in Table 12.
Appendix B Additional Experimental Results
B.1 Initial Investigation by Switching Thinking Mode
Recent works  (Zhou et al., 2025a; Zhang et al., 2025) have discovered that LRMs tend to answer harmful questions. To further validate whether the thinking trajectory influences the safety of LLMs, we compare multiple LRMs with the thinking mode switch on & off. The evaluation is performed on 2k harmful queries from Wildjailbreak  (Jiang et al., 2024a), with safety assessed using Llama-Guard-3-8B (Llama Team, 2024) as the automatic judge.
The thinking trajectory substantially undermines the safety of LRMs. As shown in Table 13, LRMs become consistently safer when the thinking mode is disabled, but their safety degrades notably once it is enabled. This trend holds across different model scales: enabling the thinking trajectory systematically increases the likelihood of harmful responses. For example, the harmful answer rate of Qwen3-32B rises by 10.55% when the mode is enabled, while the 8B and 14B variants also exhibit noticeable increases (6.04% and 3.35%, respectively). These results indicate that content generated within the thinking trajectory can directly drive LRMs to produce harmful outputs.
Table 13: Percentage of harmful answers (”Answer”) and refusals (”Refuse”) on Wildjailbreak. Higher ’Answer’ indicates lower safety. $\Delta=$ Disabled - Enabled.
Model Enabled Disabled $\Delta$ (Answer)
Answer Refuse Answer Refuse
Qwen3-8B 41.40 58.60 35.36 64.64 6.04-6.04
Qwen3-14B 35.65 64.35 32.30 67.70 3.35-3.35
Qwen3-32B 41.45 58.55 30.90 79.10 10.55-10.55
B.2 Analysis: PCA Analysis of 8B and 14B Models
We conducted a comparative analysis of models with different parameter sizes and fine-tuning methods (SafR and SafB), aiming to evaluate their impact on safety and representational clustering. The observed differences are visualized in Figure 6 and 7, while quantitative results across all configurations are reported in Table 14.
Table 14: Safety distance analysis of Qwen3 models by model size. Values represent perpendicular distances from cluster centroids to the linear decision boundary. Positive $\Delta$ values indicate movement away from the boundary (greater margin, safer region), while negative $\Delta$ values indicate movement toward the boundary (reduced margin, less safe region).
Method Qwen3-8B Qwen3-14B Qwen3-32B
Harmful Reasoning Harmful Reasoning Harmful Reasoning
$\Delta$ $\Delta$ $\Delta$ $\Delta$ $\Delta$ $\Delta$
Base 13.24 14.48 15.54 9.05 11.20 12.18
SafR 18.55 15.02 18.95 8.33 16.18 11.50
+5.30 +0.54 +3.41 -0.73 +4.98 -0.68
SafB 19.51 12.95 19.14 9.63 15.16 10.41
+6.27 -1.54 +3.60 +0.58 +3.96 -1.77
SafeChain 13.45 5.87 12.59 4.26 11.58 6.92
+0.21 -8.61 -2.95 -4.79 +0.38 -5.26
Model Scale
The 32B model consistently outperforms the 8B model across safety and clustering metrics. It exhibits a higher Safety Distance in all settings (Base, SafR, SafB), indicating better separation from harmful content. Its Silhouette Score at the Base stage (0.140) also exceeds that of the 8B model (0.120), reflecting a more structured internal representation.
Fine-Tuning Methods
Both Safety Recomposition (SafR) and Safety Backtrack (SafB) fine-tuning strategies contribute to enhanced model safety, as evidenced by the centroid shifts observed in Figure 6 and Figure 7, and further quantified by the increase in Safety Distance (e.g., for the 8B model: from 20.05 to 25.68 after SafR). Notably, SafR yields consistently superior clustering performance, with higher Silhouette Scores than SafB across both model sizes. This suggests that SafR not only improves safety alignment but also leads to more coherent and well-separated internal representations.
Figure 6: PCA of the Qwen3-8B representation space.
Figure 7: PCA of the Qwen3-14B representation space.
B.3 Preservation of Reasoning Token Usage
Table 15 reports the average token counts for GPQA and AIME benchmarks across different training methods. SafR and SafB maintain token usage close to the Base model, while improving safety. In contrast, other baselines exhibit larger deviations, suggesting potential inefficiencies in reasoning. These results demonstrate that SafR and SafB enhance model safety without significantly affecting intrinsic reasoning behavior, selectively reinforcing safety-related features while preserving core reasoning capabilities.
Table 15: Average token length on GPQA-Diamond and AIME benchmarks for Qwen3-8B, Qwen3-14B, and Qwen3-32B across different training methods.
Model Method GPQA-Diamond AIME
Qwen3-8B Base 7553.84 14895.49
SafePath 4869.04 11816.52
STAR-1 3449.49 12337.94
SafeChain 5105.35 12533.35
SafR 5212.79 14472.53
SafB 6540.14 14881.68
Qwen3-14B Base 5585.28 14419.94
SafePath 4379.39 10939.46
STAR-1 2736.43 11011.73
SafeChain 4688.32 12372.85
SafR 4303.19 13549.96
SafB 5091.59 13010.36
Qwen3-32B Base 5150.81 12634.11
SafeChain 5588.61 11198.09
SafePath 3925.96 10200.49
STAR-1 3564.03 11862.56
SafB 6452.00 12318.87
SafR 4830.90 12548.96
B.4 Analysis: Reasoning Safety Analysis
Prior work has primarily focused on improving the safety of reasoning through external interventions (safety for reasoning). In this section, we take the opposite view and explore whether a model’s internal reasoning ability can be leveraged to enhance overall safety (reasoning for safety). To evaluate our approach, we compare it against the recently proposed R2D method  (Zhu et al., 2025) as a representative baseline. The experimental results are shown in Table 16.
Table 16: Performance comparison on safety and reasoning benchmarks between our COG and the R2D baseline. Lower is better for SafB, SafR, W-JB, and JB-B; higher is better for GPQA and AIME.
Model SafB SafR W-JB JB-B GPQA AIME
Qwen3-8B 45.45 13.62 38.80 18.29 57.58 73.30
R2D 19.32 0.96 12.80 84.15 41.92 47.92
SafR(Ours) 13.41 2.86 9.20 71.95 56.82 76.25
SafB(Ours) 17.05 2.05 8.00 73.17 54.30 77.50
Based on the experimental results shown in the Table 16, we observe that applying reasoning-for-safety methods to LRMs indeed improves safety performance. However, similar to SafeKey and Star-1, these methods tend to sacrifice reasoning ability, failing to strike a proper balance between safety and reasoning capabilities. Although our methods partially address this issue, maintaining the integrity of reasoning ability remains a significant challenge.
Appendix C Self-Jailbreak Category Example
C.1 Benign Reframing
The example of Benign Reframing in Self-Jailbreak is presented in Figure 8, where red text represents their Self-Jailbreak act.
C.2 Logical Fallacies
The example of Logical Fallacies in Self-Jailbreak is presented in Figure 9, where red text represents their Self-Jailbreak act.
C.3 Warning
The example of Warning in Self-Jailbreak is presented in Figure 10, where red text represents their Self-Jailbreak act.
C.4 Harm Identification
The example of Warning in Harm Identification is presented in Figure 11, where LRM answers the harmful quesiton as the regular one.
Appendix D Prompt Design
In this section, we present the prompt designs used in our COG framework and other experiments.
D.1 Extraction Prompt
This subsection provides both the base prompts and few-shot examples used in the extraction task.
Prompt
The basic prompt template used for extraction is shown in Figure 12.
Few-Shot
We also include few-shot examples to guide the model during extraction, as illustrated in Figure 13.
D.2 Classification Prompt
Prompt
The prompt used for the classification task is outlined below in Figure 14.
D.3 Safety Recomposition Prompt
For the Safety Recomposition stage, our prompt is constructed by concatenating several components. Specifically, based on the classification result, we extract a corresponding rewrite instruction from the “sub_prompts” field in a JSON file. Then, we combine the ”main_prompt”, the selected ”sub_prompt”, and a ”format_prompt” to form the final prompt. The detailed structure is as shown in Figure 15
D.4 Safety Backtrack Prompt
In the Safety Backtrack stage, we further extend the prompt structure from the Safety Recomposition process. In addition to the previous components, we incorporate a transition phrase from “contextual_transition_phrases”, selected based on the classification result. This helps guide the model more smoothly and maintain coherence in the final output. The detailed structure is as shown in Figure 16
D.5 Integration Prompt
This stage integrates the outputs from previous modules into a unified final response. The detailed prompt is as shown in Figure 17
D.6 BoN Ranking Prompt
In this step, we employ a Best-of-N (BoN) ranking mechanism to select the most appropriate final output from multiple candidates generated by the model. The ranking prompt used to guide this selection process is shown in Figure 18.
Figure 8: Example of Benign Reframing.
Figure 9: Example of Warning.
Figure 10: Example of Warning.
Figure 11: Example of Harm Identification.
Figure 12: Example of the basic extraction prompt used in the extraction stage.
Figure 13: Few-shot examples used in the extraction prompt during the extraction stage.
Figure 14: Example of the basic classification prompt used in the classification stage.
Figure 15: Prompt structure used in the Safety Recomposition stage, formed by concatenating the main prompt, selected sub-prompt based on classification results, and format prompt.
Figure 16: Prompt structure used in the Safety Backtrack stage, incorporating contextual transition phrases based on classification results.
Figure 17: Prompt structure used in the Safety Backtrack stage, incorporating contextual transition phrases based on classification results.
Figure 18: Prompt structure used in the Best-of-N (BoN) ranking stage to select the best output from multiple candidates.