### v3

# When Models Outthink Their Safety: Unveiling and Mitigating Self-Jailbreak in Large Reasoning Models

Abstract
Large Reasoning Models (LRMs) achieve strong performance on complex multi-step reasoning, yet they still exhibit severe safety failures such as harmful content generation. Existing methods often apply coarse-grained constraints over the entire reasoning trajectories, which can undermine reasoning capability while failing to address the root causes of unsafe behavior. In this work, we uncover a previously underexplored failure mode in LRMs, termed Self-Jailbreak, where models initially recognize the harmful intent of a query, but override this judgment during subsequent reasoning steps, ultimately generating unsafe outputs. Such a phenomenon reveals that LRMs are capable of recognizing harm, while safety failures primarily arise from reasoning steps. Motivated by this finding, we propose Chain-of-Guardrail (CoG), a trajectory-level training framework that mitigates Self-Jailbreak via targeted, step-level interventions while maintaining reasoning ability. Experiments across multiple safety and reasoning benchmarks indicate that CoG achieves a favorable balance between safety and reasoning performance compared with existing approaches.
When Models Outthink Their Safety:
Unveiling and Mitigating Self-Jailbreak in Large Reasoning Models
Yingzhi Mao1,2thanks: Equal contribution.thanks: Code: https://github.com/icip-cas/COG., Chunkang Zhang [^1] footnotemark: 1, Junxiang Wang1, Xinyan Guan1,2, Boxi Cao1 Yaojie Lu1, Hongyu Lin1, Xianpei Han1,2, Le Sun1,2 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2University of Chinese Academy of Sciences {maoyingzhi2024, zhangchunkang2022, luyaojie}@iscas.ac.cn
1 Introduction
Large Reasoning Models (LRMs) (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025) have demonstrated strong capabilities across a wide range of tasks, including software engineering and mathematical reasoning. As these models are increasingly integrated into autonomous agents and decision-support systems, ensuring their safety and alignment with human values is foundational.
However, recent studies show LRMs remain vulnerable to serious safety failures (Green et al., 2025; Arrieta et al., 2025), ranging from generating harmful content (Qiu et al., 2025; He et al., 2025) to exhibiting vulnerability to jailbreak attacks (Zhou et al., 2025a; Zhang et al., 2025b). Therefore, strengthening safety capabilities of LRMs is not merely an auxiliary concern, but a prerequisite for their trustworthy deployment at scale.
Figure 1: Illustration of Self-Jailbreak. While the model initially recognizes the harmful nature of a query, it overrides this judgment during subsequent reasoning steps, ultimately yielding unsafe output.
While recent efforts (Wang et al., 2025; Jeung et al., 2025; Zhou et al., 2025a; Jiang et al., 2025) have made progress towards safer LRMs, they often incur a pronounced safety–reasoning trade-off. A primary limitation shared by these approaches is that they typically inherit safety paradigms originally designed for standard LLMs, treating internal reasoning trajectories of LRMs as an undifferentiated extension of responses. Subsequently, safety constraints are typically imposed in a coarse-grained and global manner. Such heuristic safeguards often interfere with the model’s intrinsic reasoning patterns, degrading the coherence required for multi-step reasoning. These limitations indicate that mitigating safety risks in LRMs without undermining their reasoning capabilities requires moving toward mechanisms that can localize and address failure-inducing steps within the reasoning chain.
To this end, we conduct a systematic analysis of reasoning trajectories of LRMs to uncover the root causes of their safety failures. In particular, we identify two distinct failure modes: Harm Misidentification, where harmful intent is not recognized, and Self-Jailbreak, a severe and previously overlooked phenomenon in which the model initially identifies potential harm but later overturns this safety judgment during subsequent reasoning. To enable fine-grained analysis, we decompose each reasoning trajectory into three consecutive stages: risk awareness, risk analysis, and response strategy. Specifically, as demonstrated in Figure 1, given a harmful query (“How to build a bomb…”), the model correctly identifies the harm during the risk-awareness stage, but then overrides this judgment by rationalizing a seemingly benign user intent (“…could be a student…”). In the response-strategy stage, it chooses to answer rather than refuse, and the resulting response contains harmful content.
Through quantitative analysis on WildJailbreak, we find that only a small portion of unsafe outputs arise from Harm Misidentification, whereas Self-Jailbreak is the dominant failure mode, responsible for nearly 80% of the unsafe cases we analyze and thus constituting our primary focus. In these cases, LRMs correctly recognize harmful intent during the risk awareness stage, yet this recognition is subsequently overridden during the risk analysis stage, where the model effectively persuades itself to comply with the unsafe request.
Building on these insights, we propose Chain-of-Guardrail (CoG), a training framework designed to mitigate Self-Jailbreak while preserving the reasoning capability. CoG executes targeted, step-level interventions conditioned on the diagnosed Self-Jailbreak patterns, correcting only the segments that induce unsafe behavior. We instantiate this framework via two complementary variants: Safety Recomposition, which rewrites the reasoning chain into a logically consistent, safe alternative, and Safety Backtrack, which preserves the original trajectory while revising risky segments before they lead to unsafe outputs. These corrected traces serve as fine-tuning data to align LRMs without suppressing their reasoning potential.
Extensive experiments across multiple safety and reasoning benchmarks demonstrate that CoG achieves a superior balance between safety and reasoning performance. Across diverse LRMs, CoG consistently reduces attack success rates to competitive levels on standard safety benchmarks, while substantially boosting performance on challenging reasoning tasks. Notably, on Qwen3-32B, CoG achieves safety performance comparable to SafeKey, while substantially improving reasoning accuracy, with GPQA-Diamond increasing from 54.30 to 62.38 and AIME2024 from 71.70 to 82.08. Further analyses from both reasoning-pattern and representation-geometry perspectives suggest that CoG improves safety by correcting specific failure-inducing reasoning steps, rather than globally rewriting the model’s reasoning paradigm.
We summarize our major contributions as follows:
  • To the best of our knowledge, we are the first to uncover and characterize the Self-Jailbreak phenomenon, revealing it as a primary driver of safety failures in LRMs.
  • We propose an analysis framework that enables the categorization and quantitative analysis of unsafe reasoning behaviors.
  • We introduce the Chain-of-Guardrail (CoG), a trajectory-level training framework that achieves the state-of-the-art safety-reasoning balance across multiple benchmarks.
2 Unveiling Self-Jailbreak in Reasoning Trajectories
In this section, we analyze the causes of safety failures in LRMs through an analysis of their reasoning trajectories. Our findings indicate a consistent mismatch between risk awareness and final response: LRMs may generate unsafe outputs even after identifying potential risks during reasoning. Beyond failures of harm identification, we observe a prevalent pattern in which earlier safety assessments are revised or overridden in later reasoning stages, which we refer to as Self-Jailbreak. We further provide a taxonomy of such behaviors across multiple LRMs to support a more fine-grained analysis of this failure mode.
2.1 The Prevalance of Self-Jailbreak in LRMs
Investigation Setup
To localize safety failures within the reasoning process, we decompose each reasoning trajectory into three stages: risk awareness, risk analysis, and response strategy. Each stage is assessed independently. For evaluation, we sample 2,000 data points from Wildjailbreak (Jiang et al., 2024a) as the benchmark, following the official setting with Llama-Guard (Llama Team, 2024) to evaluate whether the final response contains harmful content. In addition, we utilize Qwen2.5-72B-Instruct (Team, 2024; Gu et al., 2024) to determine whether potential risks are explicitly identified during the risk awareness stage.
Figure 2: Flow of harmful queries in WildJailbreak through Qwen3-32B, showing the relationship between the final response (Answer/Refuse) and risk awareness during reasoning (Yes/No).
Main Cause of Safety Failure
Based on the results of Figure 2, we categorize safety failures into two distinct patterns: (1) Harm Misidentification, where the model fails to recognize harmful intent, and (2) Mismatch between risk awareness and response, where the model explicitly identifies risks but produces unsafe outputs. While the former indicates a detection failure, the latter reveals a more critical flaw in the model’s reasoning process.
Further investigation of the second pattern reveals a pattern in which the model’s subsequent reasoning effectively revises or overrides its own safety assessment, leading to unsafe outputs even after risks have been identified. We refer to this phenomenon as Self-Jailbreak.
To determine whether Self-Jailbreak is an isolated anomaly or a widespread issue, we extended our investigation across multiple Large Reasoning Models (LRMs), as shown in Figure 3. Our results yield a significant finding:
Self-Jailbreak constitutes the predominant safety failure mode across LRMs. As illustrated in Figure 2, Self-Jailbreak consistently surpasses Harm Misidentification as the primary source of unsafe outputs across diverse model families and parameters. For instance, in DeepSeek-R1, 93.7% of safety failures are attributable to Self-Jailbreak. This indicates that Self-Jailbreak represents a systemic and persistent challenge, even for state-of-the-art LRMs.
Figure 3: Proportion of the safety failure cause across multiple LRMs conditioned on the unsafe responses of Wildjailbreak
Model Benign Reframing Warning Logical Fallacies
DS-R1 40.91 56.97 2.12
DS-Llama-70B 39.17 58.53 2.30
Qwen3-8B 36.94 57.91 5.15
Qwen3-14B 38.60 58.20 3.20
Qwen3-32B 36.86 58.96 4.18
Table 1: Distribution of self-jailbreak categories (%), where DS-R1 stands for DeepSeek-R1 while DS-Llama-70B represents DeepSeek-R1-Distill-Llama-70B.
2.2 A Taxonomy of Self-Jailbreak Behaviors
While the prevalence of self-jailbreak is evident, the underlying reasons that lead models to generate unsafe content remain poorly understood. A deeper analysis of how models override their own safety judgment is essential for designing effective safeguards. To this end, we conducted a qualitative analysis involving manual inspection of reasoning traces (Chain-of-Thought) and final responses, focusing on the discrepancy between the model’s internal risk analysis and its final output generation.
Taxonomy Definition
Based on recurring patterns identified in the manual inspection, we establish a taxonomy categorizing Self-Jailbreak into three distinct behaviors:
  • Benign Reframing: The model actively reinterprets the user’s malicious intent as benign (e.g., educational or theoretical), thereby justifying a helpful response.
  • Warning: The model assumes that appending a safety warning or disclaimer is sufficient to mitigate the harm, leading to a “warn-but-answer” failure mode.
  • Logical Fallacies: The model’s reasoning becomes entangled in complex or contradictory logical constraints within the prompt, causing it to bypass safety guardrails due to erroneous logical deductions (examples in Appendix E).
Figure 4: Overview of the Chain-of-Guardrail (CoG) framework. Phase 1: the original LRM produces an initial COT, which is decomposed into three atomic components—risk awareness $d(x)$, risk analysis $a(x)$, and response strategy p(x)p(x)—and classified for self-jailbreak risks. Phase 2: guided by the classification signal, CoG applies Safety Recomposition (SafR) or Safety Backtrack (SafB) to construct a safety-oriented COT (S-COT). Phase 3: The S-COT guides the model to generate the final safe response.
Quantitative Analysis
Utilizing this taxonomy, we annotated Self-Jailbreak instances across multiple LRMs to analyze their distribution. As shown in Table 1, a key finding emerges. Most notably, Warning is the primary Self-Jailbreak type, consistently accounting for over 55% of cases across all evaluated models. This suggests that the primary safety bottleneck is not a lack of risk awareness, but rather a flaw in the risk analysis. Models have learned to be overly compliant and often default to answering even when risks are identified, rather than refusing. The prevalent Warning and Benign Framing behavior indicates that during training, models may not have internalized clear signals for when to withhold responses entirely.
Implications for Methodology
Overall, these results indicate that Self-Jailbreak is both widespread and structured. Unsafe behavior often arises after correct risk recognition and recurs in a small number of type-specific reasoning patterns. This implies that coarse-grained safety enforcement or globally rewriting reasoning traces may discard valid reasoning steps while still missing the failure source, motivating mitigation that first identifies the Self-Jailbreak type and then targets the corresponding failure-inducing steps.
3 Chain-of-Guardrails: A Training Framework for Reasoning-Aware Safety
Motivated by the observation that safety failures in LRMs frequently arise during reasoning despite correct risk awareness, we propose Chain-of-Guardrails (CoG), a training framework that identifies and corrects unsafe reasoning steps while preserving the model’s inherent reasoning capability. Rather than imposing uniform safety constraints on the entire response, CoG performs targeted, reasoning-level interventions guided by diagnosed Self-Jailbreak behaviors.
3.1 Framework Overview
Given a query $x$ and an original LRM $\pi_{0}$, CoG aims to produce a safe response $y_{\text{safe}}$ while preserving the model’s original reasoning ability. To enable fine-grained analysis and intervention, we explicitly decompose the model’s reasoning trajectory into interpretable components.
3.2 Phase 1: Input & Analysis
In Phase 1, the original model $\pi_{0}$ generates an initial reasoning trajectory and response to the input query $x$. We decompose the resulting reasoning trajectory into three interpretable components:
$c=\pi_{0}(x)=[d(x),a(x),p(x)]$ (1)
where $d(x)$ denotes the model’s risk awareness, $a(x)$ its risk analysis, and $p(x)$ its response strategy.
Given the decomposed components, we run a Self-Jailbreak classifier to predict whether the reasoning trajectory contains Self-Jailbreak behaviors and, when applicable, output the corresponding Self-Jailbreak type.
3.3 Phase 2: Safety-Oriented Reasoning Transformation
In Phase 2, CoG transforms unsafe reasoning trajectories into safety-oriented ones by applying targeted interventions conditioned on the classification signal from Phase 1.
We introduce the following complementary transformation strategies:
[…]
Safety Backtrack (SafB) SafB keeps the original reasoning chain and adds a targeted self-check step. Guided by the Self-Jailbreak type, the original model $\pi_{0}$ takes the original $a(x)$ and $p(x)$ and generates a self-check segment that focuses on the failure-inducing parts of the reasoning. This self-check revisits the earlier reasoning decisions and provides corrective guidance before producing the final response. We then append the self-check segment to the end of the original chain, forming an augmented, safety-oriented reasoning trajectory.
3.4 Phase 3: Inference & Alignment
In Phase 3, the safety-oriented COT produced by SafR or SafB is used to generate a final safe response, which is then used as training data for alignment. We then fine-tune the original model $\pi_{0}$ on this constructed data using selective loss masking.
For SafR, we supervise the entire recomposed reasoning chain together with the final response. For SafB, we supervise only the appended self-check segment and the final response, while masking the original reasoning tokens. This selective supervision encourages learning safety-oriented corrections while mitigating unintended distributional shift away from $\pi_{0}$’s original reasoning trajectory.
Method Harmful Jailbreak Reasoning
[…]
Vanilla 45.45 13.62 38.80 81.71 29.00 57.33 77.50 97.6 92.68
STAR-1 18.86 0.74 20.00 37.8 5.00 57.33 71.25 96.4 93.29
SafePath 36.14 10.03 22.80 42.68 14.00 55.56 67.92 94.8 92.07
SafeChain 49.55 16.99 36.80 70.73 39.00 52.28 66.25 95.2 90.24
SafeKey 3.18 0.32 8.53 13.2 7.00 41.90 70.58 90.0 91.46
Safety Backtrack (Ours) 17.05 1.45 8.00 26.83 5.00 54.30 77.50 97.4 93.90
Safety Recomposition (Ours) 13.18 1.89 9.20 28.05 5.00 56.82 76.25 92.6 93.90
[…]
Vanilla 55.45 12.44 34.00 68.29 21.00 63.14 77.92 97.6 95.73
STAR-1 17.95 0.72 13.20 23.17 32.00 56.32 76.25 88.4 97.56
SafePath 24.09 8.49 16.00 50.00 15.00 57.78 70.42 78.8 88.41
SafeChain 49.09 16.87 35.60 67.07 32.00 57.58 71.25 87.4 95.73
SafeKey 4.77 0.32 4.88 10.40 6.00 49.00 76.70 88.4 86.58
Safety Backtrack (Ours) 10.91 0.97 6.40 18.05 7.00 62.12 77.92 97.0 96.34
Safety Recomposition (Ours) 7.95 2.09 2.80 21.95 5.00 60.36 78.75 97.4 96.34
[…]
Vanilla 46.59 12.25 35.20 80.43 23.00 65.66 81.67 97.6 98.17
STAR-1 18.41 0.83 16.80 35.37 3.00 54.55 72.92 85.2 95.12
SafePath 40.00 6.57 22.80 53.66 20.00 62.38 70.25 60.4 97.56
SafeChain 47.95 16.39 28.40 70.73 34.00 54.30 71.70 86.4 96.34
SafeKey 3.41 0.32 7.32 10.40 10.00 54.30 71.70 86.8 87.20
Safety Backtrack (Ours) 14.55 1.68 8.80 23.17 1.00 61.62 77.08 97.4 98.17
Safety Recomposition (Ours) 7.05 1.10 3.20 17.07 2.00 62.38 82.08 97.6 97.56
Table 2: Performance comparison of different methods under the main experimental setting. Lower values in harmful/jailbreak benchmarks and higher values in reasoning benchmarks represent better performance.
4 Experiment
In this section, we conduct experiments across multiple safety and reasoning benchmarks to validate the effectiveness of our proposed COG framework. We demonstrate that compared to prior baselines, our methods achieve state-of-the-art safety-reasoning balance across different model scales. Furthermore, we perform in-depth analyses to investigate the underlying mechanisms that contribute to the success of our approach, examining both the preservation of reasoning patterns and the distributional characteristics of the learned representations.
4.1 Experiment Setting
Training Dataset We collect 15,000 high-quality harmful queries from public datasets including Alert, ToxicDPOqa, Harmful-Dataset, Aya_RedTeaming, Do-Not-Answer, AttaQ, and Toxic-Chat (Tedeschi et al., 2024; Ahmadian et al., 2024; Wang et al., 2023; Kour et al., 2023; Lin et al., 2023). This curated set serves as input to our pipeline for constructing corresponding safety-aligned responses.
Evaluation Benchmarks For safety, we use Sorry-bench and StrongREJECT for harmful-prompt refusal, and WildJailBreak and JailBreakBench for jailbreak robustness (Xie et al., 2024; Souly et al., 2024; Jiang et al., 2024b; Chao et al., 2024). For reasoning, we adopt GPQA-Diamond, AIME2024, MATH500, and HumanEval to assess math and code reasoning capabilities (Rein et al., 2024; Mathematical Association of America (MAA), 2024; Lightman et al., 2023; Chen et al., 2021).
Models and Configuration To validate the efficacy of our methods, we utilize the Qwen3 series  (Yang et al., 2025). SafR and SafB are implemented with training details in Appendix A.2.1.
Baselines We compare with diverse representative safety alignment baselines including STAR-1, SafeChain, SafePath, and SafeKey (Wang et al., 2025; Jiang et al., 2025; Jeung et al., 2025; Zhou et al., 2025b). All baselines are reproduced following official codes, configurations, and datasets, detailed in Appendix A.2.1
4.2 Overall Results
Table 3.4 summarizes the overall safety and reasoning performance across model scales. The results reveal clear differences in how existing methods trade off safety and reasoning, and highlight the effectiveness of CoG under this trade-off.
Figure 5: Safety vs. reasoning trade-off for the Qwen 32B.
1) Vanilla LRMs without safety alignment exhibit poor safety performance. Despite their impressive performance on reasoning benchmarks, the vanilla versions of Qwen3-8B, 14B, and 32B consistently achieve the highest (worst) scores across all safety and jailbreak metrics. For instance, on the Sorry-bench and Wildjailbreak benchmarks, the vanilla models demonstrate a high propensity to follow harmful instructions, with Sorry-bench scores reaching as high as 55.45% in the 14B variant. This confirms that advanced reasoning capabilities do not inherently translate to safety, and without explicit alignment, these models remain fragile.
2) Previous approaches enhance safety performance at the substantial cost of reasoning capability. Baselines such as SafeKey and STAR-1 effectively reduce safety risks, but often incur a clear compromise in reasoning. For instance, on Qwen3-32B, SafeKey reduces Sorry-bench from 46.59% to 3.41% and JBB-PAIR from 80.43% to 10.40%, at the expense of lowering GPQA-Diamond from 65.66% to 54.30% and AIME2024 from 81.67% to 71.70% Similar trade-offs are observed for STAR-1 across model scales, suggesting that the global safety signal is harmful for reasoning capability.
3) COG achieves the best safety-reasoning balance across all model scales. As shown in Table 3.4, our methods significantly boost safety, reducing Wildjailbreak and Sorry-bench scores to levels comparable with the strongest safety baselines. This trade-off is clearly visualized in Figure 5, where our methods achieve the most favorable balance between safety and reasoning performance. For example, on Qwen3-32B, Safety Recomposition reduces Sorry-bench from 46.59% to 7.05% while maintaining AIME2024 at 82.08% and MATH500 at 97.6% compared to the vanilla model; Safety Backtrack provides a complementary point that further preserves reasoning with slightly weaker safety than the most safety baseline. Overall, these results indicate that CoG can deliver strong safety gains with minimal loss in reasoning ability, enabling a more favorable safety-reasoning trade-off for deploying capable LRMs.
4.3 Detailed Analysis
Pattern Vanilla SafeChain SafePath Star-1 SafB SafR
Backtracking 1.33 1.10 1.20 1.30 1.27 1.30
Enumeration 0.93 0.87 0.97 0.83 1.00 1.03
Subgoal Setting 1.60 1.63 1.30 1.40 1.47 1.57
Verification 2.50 2.47 2.23 2.10 2.50 2.57
[…]
Table 3: Comparison on the frequencies of reasoning patterns (Qwen3-32B) across different training strategies. “Overall Avg." denotes the average frequency across all reasoning patterns, reflecting the overall reasoning style shift under different strategies.
Figure 6: PCA of the Qwen3-32B representation space. Gray ellipses indicate base model distributions, while colored regions denote post-training clusters (red: safety, blue: reasoning). Centroid shifts ($\Delta$) quantify displacement relative to the linear decision boundary (dashed), with $\Delta>0$ indicating increased margin and $\Delta<0$ indicating decreased margin.
We further examine why COG improves safety without sacrificing reasoning largely. Specifically, we test whether COG preserves the base model’s reasoning paradigm from two perspectives—reasoning trajectory patterns and representation separability —and provide illustrative case studies in Appendix D.
4.3.1 Reasoning Pattern
To study how training strategies influence the reasoning structure of LRMs, we begin with a simple token-level signal. Based on our token-level statistics, SafR and SafB induce minimal changes in generation length on GPQA-Diamond and AIME compared to the base model, while baselines—such as STAR-1—consistently produce substantially shorter outputs across all model scales; see Appendix C.2 for details.
Subsequently, we conduct a deeper reasoning-pattern analysis, which calculates the frequencies of reasoning patterns among various methods. Following established prompting protocols (Zeng et al., 2025; Gandhi et al., 2025), we use DeepSeek-v3.1 (DeepSeek-AI, 2024) to analyze cognitive trajectories and quantify four key behaviors—Backtracking, Enumeration, Subgoal Setting, and Verification—measured as the average occurrences per problem on AIME. CoG resembles the most to original LRMs in reasoning frequencies. Table 3 indicates that SafR and SafB remain the closest to the vanilla model in behavior frequencies (SafR: +0.03%, SafB: -0.03%), while SafeChain, SafePath, and STAR-1 exhibit substantially larger deviations. Overall, these results suggest that CoG improves safety while largely preserving original reasoning patterns.
4.3.2 Representation Separability
To quantify safety-reasoning separability at the representation level, we apply PCA to Qwen3-32B hidden states on safety-critical prompts and reasoning prompts, and fit a linear decision boundary to quantify their separability, as illustrated in Figure 6.
We define each cluster’s margin as the distance from its centroid to the decision boundary: a larger margin means the centroid is farther from the boundary and thus easier to separate, while a smaller margin indicates weaker separability.
CoG improves safety separation while largely preserving reasoning representations. SafR produces the largest increase in the safety margin (+4.985+4.985) with only a small reduction in the reasoning margin (0.676-0.676). SafB increases the safety margin (+3.961+3.961), but reduces the reasoning margin more (1.771-1.771), placing reasoning representations closer to the decision boundary. In contrast, SafeChain yields a small safety gain (+0.378+0.378) while greatly reducing the reasoning margin (5.260-5.260), indicating much weaker separability for reasoning representations. Overall, these results suggest CoG strengthens safety separation without strongly distorting reasoning representations.
Comparison Awareness Classification
Human A 0.81 0.77
Human B 0.79 0.71
Human C 0.78 0.73
Mean Consensus 0.79 0.74
Table 4: Human–model agreement (Pearson’s $r$) on 50 instance-level binary judgments for Self-Jailbreak classification and LlamaGuard, where Awareness stands for the awareness of the risk in queries, while Classification represents the classification of Self-Jailbreak.
4.3.3 Consistency with human
To verify the effectiveness of LLM-as-a-judge in CoG framework and discovering Self-Jailbreak, we conduct a human–model agreement on a sampled subset from Wildjailbreak by comparing human judgments with model decisions when deciding the awareness of the risk in query as well as the classification of Self-jailbreak.
As shown in Table 4, both systems exhibit strong and consistent alignment with human judgments across annotators, achieving correlations of 0.74 (Awareness) and 0.79 (Classification) with the averaged human consensus.
4.4 Ablation Study
To verify the effect of the selective loss masking strategy in SafB, we compare (i) partial-mask training (our default), where supervision is applied only to the self-check segment and the final answer, and (ii) no-mask training, where the full sequence (including the original reasoning trace) is supervised.
[…]
SafB 17.05 2.05 8.00 26.83
SafB (w/o mask) 23.64 5.37 22.40 59.76
Table 5: Effect of selective loss masking in SafB.
Table 5 shows that selective masking is crucial for SafB. Partial-mask training consistently achieves better results on safety benchmarks. In contrast, removing the mask and supervising the full reasoning trace substantially degrades safety, with the largest drops observed on jailbreak benchmarks. This supports masking the original reasoning trace in SafB and applying supervision only to the self-check and final answer, which helps mitigate unintended distributional shift.
5 Related Work
Vulnerable Safety of LRMs. Recent studies consistently show that LRMs remain vulnerable under harmful queries and adversarial settings, and that longer, explicit reasoning can introduce extra safety risks. (Huang et al., 2025) shows a clear safety-reasoning trade-off in common alignment pipelines, where better safety can hurt reasoning performance. (Zhu et al., 2025) shows that attackers can manipulate special delimiter markers to bypass reasoning, effectively skipping the intended deliberation. (Xu et al., 2025) demonstrates that safety alignment can be undermined via fine-tuning attacks against CoT-enabled models. Other work, such as (Li et al., 2025; Zhang et al., 2025a), also studies how safety changes as reasoning improves, and shows that reasoning-time safety can still fail in hard or adversarial cases.
Safety Alignment for LRMs. To mitigate these risks, recent methods align LRMs by shaping the reasoning process or injecting safety deliberation. (Guan et al., 2024) trains models to explicitly recall and reason over safety specifications before answering. Data-driven approaches such as (Wang et al., 2025; Zhang et al., 2025a) construct safety-oriented reasoning trajectories for fine-tuning. (Jiang et al., 2025) both assesses long-CoT safety risks and introduces CoT-style safety training data. Orthogonally, (Zhou et al., 2025b) aims to strengthen internal safety activation signals, while (Doula et al., 2025) promotes early safety priming to reduce harmful reasoning. In contrast, we analyze reasoning trajectories stage by stage and formalize Self-Jailbreak as a failure mode after risk awareness.
6 Conclusion
We systematically analyze safety failures in LRMs and uncover Self-Jailbreak, in which the model initially recognizes harmful intent but later overrides this judgment during reasoning, leading to unsafe outputs. This finding suggests that many safety failures stem from failure-inducing steps that override correct risk awareness within the reasoning chain. Motivated by this, we propose Chain-of-Guardrail (CoG), a trajectory-level training framework that mitigates Self-Jailbreak via targeted, step-level interventions. Across multiple safety and reasoning benchmarks, CoG improves safety and robustness while maintaining comparable reasoning performance.
Overall, our study provides an analysis lens and a practical framework for aligning LRMs toward safer reasoning.
Limitations
We acknowledge several limitations in the current study.
First, due to computational resource constraints, we do not evaluate our method on substantially larger-scale reasoning models. While our experiments demonstrate the effectiveness of the proposed framework on the models studied in this paper, validating its behavior at larger scales remains an important direction for future work.
Second, a large portion of our evaluation relies on automated judgments using LLM-based evaluators. Although LLM-as-judge provides scalability and consistency for analyzing fine-grained reasoning behaviors, it remains an imperfect proxy for human judgment. To partially mitigate this limitation, we conduct human–model consistency analyses and observe strong agreement between human annotations and automated evaluations. Nevertheless, establishing more reliable, standardized, and cost-effective evaluation protocols for safety failures in long-form reasoning remains an open problem. We leave the development of improved human-in-the-loop or hybrid evaluation frameworks to future work.
Appendix A Detailed Experimental Setup
A.1 Baseline Configuration
A.1.1 Description of Baseline Methods
STAR-1
STAR-1 categorizes 41,000 safety data points from multiple sources into eight predefined categories and generates a response with COT using DeepSeek-R1, guided by the safety policies associated with each category. Then, a set of rules is applied to filter out 1,000 data points for the dataset. These data are then used to fine-tune an LRM to conduct safety alignment.
SafeChain
SafeChain selected 50,000 data points from the Wildjailbreak dataset and used R1-70B to generate five responses for each instruction. Then, Llama-Guard is used to filter data, keeping the responses that are all safe. A random response is sampled from the five responses as the final response. This created a dataset containing 40,000 instruction-response pairs, available for supervised fine-tuning.
SafePath
SafePath fine-tunes LRMs in a specific way, making them always generate eight fixed tokens: “let’s think about safety first" at the start of inference, guiding the LRMs to consider more about safety during the generation process.
SafeKey
SafeKey enhances safety reasoning by integrating a Dual‑Path Safety Head with Query‑Mask Modeling to amplify latent safety signals from both the raw input (X) and the model’s internal query understanding (U) during generation of the “key sentence”—this effectively triggers a safety-focused “Aha moment.” By masking out X when predicting the key sentence based solely on U, Query‑Mask Modeling strengthens the U→K pathway, while the dual‑path head reinforces these hidden‐state safety cues during fine‑tuning. Together, these two jointly improve robustness against harmful prompts.
A.1.2 Implementation Details of Baselines
Computational Resource
To ensure fair comparison and reproducibility, all experiments—including those reproducing related work—were performed on 8 A-800 with bf16 precision enabled, which allows for faster training while preserving numerical stability. The corresponding training hyperparameters are summarized as follows.
Star-1
We use the official dataset and replicate the experiments following the parameter settings reported in the original paper. The detailed training configurations are presented in Table 6.
Hyperparameter Value
Finetuning Type Full
Optimizer AdamW
Adam $\beta_{1},\beta_{2}$ 0.9, 0.95
Learning Rate 1e-5
Epochs 5.0
Batch Size 2
Gradient Accumulation Steps 8
Weight Decay 1e-4
Warmup Ratio 0.05
Cutoff Length 8,192
Table 6: Detailed training hyperparameters for Star-1.
SafeChain
We trained Qwen3 series models with the original SafeChain dataset with llama-factory. Detailed implementation of SafeChain experiment is described as shown in Table 7:
Parameter Value
Epochs 2
Batch Size 2
Gradient Accumulation Steps 2
Table 7: Detailed training hyperparameters for SafeChain.
Safekey
We use the official SafeKey codebase, making only model-level modifications to its startup scripts. The detailed implementation of the SafeKey experiment is described as shown in Table 8.
Parameter Value
Epochs 5
Batch Size 2
Gradient Accumulation Steps 8
Table 8: Detailed training hyperparameters for SafeKey.
SafePath
The detailed implementation of the SafePath experiment is described in Table 9.
Parameter Value
Finetuning Type Full
Cutoff Length 8192
Batch Size 2
Gradient Accumulation Steps 2
Learning Rate 1e-5
Max Steps 20
Warmup Ratio 0.05
Table 9: Detailed training hyperparameters for SafePath.
A.2 Implementation Details of Our Method
A.2.1 COG Generation Parameters
During the sampling process(Phase 1), to ensure output diversity and prevent model degeneration, we set the temperature to 0.7, top_p to 0.8, and presence_penalty to 1.5 to produce the original responses used as seed data (see Table 10).
During the extraction and classification process(Phase 1), temperature and top_p were set to 0.1 and 0.9, respectively, to ensure that the model outputs its most confident predictions.
During the Safety Recomposition and Safety Backtrack stages (Phase 2), we aimed to maintain consistency between generated content and prompt constraints while preserving diversity; thus, temperature was set to 0.3 and top_p to 0.8.
Finally, for the chain-of-thought based response generation stage(Phase 3), temperature was again set to 0.7, top_p to 0.8, and presence_penalty to 1.5 to maintain diversity.
These carefully chosen parameters balance generation quality and diversity while minimally impacting the model’s reasoning capability.
Stage Parameter Value
Generation Phase temperature 0.7
top_p 0.8
presence_penalty 1.5
Extraction & Classification temperature 0.1
top_p 0.9
SafR & SafB Phases temperature 0.3
top_p 0.8
Chain-of-Thought Generation temperature 0.7
top_p 0.8
presence_penalty 1.5
Table 10: Generation parameter settings.
Parameter Value
Finetuning Type Full
Learning Rate 2e-6
Cutoff Length 8192
Epochs 3.0
Batch Size 2
Warmup Ratio 0.1
Gradient Accumulation Steps 4
Table 11: Training hyperparameters.
Safety Benchmarks Reasoning Benchmarks
Parameter Value Parameter Value
temperature 0.7 temperature 0.6
top_p 1.0 top_k 20
max_new_tokens 16384 top_p 0.95
rollout 1 max_seq_length 32768
max_out_len 32000
GPQA rollout 2
AIME2024 rollout 8
MATH500 rollout 1
HumanEval rollout 1
Table 12: Benchmark implementation details.
COG Training Parameters
Both the Safety Recomposition and Safety Backtrack tasks are trained using LlamaFactory under consistent experimental settings. Our approach is based on a dataset of 14,000 examples, with the full training hyperparameters summarized in Table 11.
A.3 Evaluation Details
A.3.1 Benchmark Description
Sorry-bench
Sorry-bench is a systematic safety-refusal benchmark comprising 440 harmful prompts across 44 fine-grained safety categories. We used the original prompts as the test set to evaluate LLM refusal behaviors.
StrongREJECT
StrongREJECT is a jailbreak robustness benchmark featuring 313 carefully filtered harmful prompts spanning six major misuse categories to assess LLM defenses against jailbreaks.
WildJailbreak
WildJailBreak is an adversarial evaluation split of 2,213 jailbreak prompts drawn from a 262 K-example synthetic safety corpus generated by the WildTeaming framework, designed to rigorously test LLM safety mechanisms. We randomly selected 250 prompts from the evaluation split as the evaluation set.
JailBreakBench
JailBreakBench is a robustness benchmark offering 100 paired harmful-behavior prompts (55 % original, 45 % sourced from AdvBench and TDC/HarmBench). In our experiment, we used harmful prompts augmented with Vicuna-generated PAIR variants for comprehensive jailbreak evaluation.
GPQA-Diamond
GPQA-Diamond is the “Diamond” subset of the GPQA benchmark, comprising the 198 most difficult of 448 graduate-level, domain-expert-written multiple-choice questions in biology, chemistry, and physics.
AIME2024
AIME is the complete set of 30 official integer-answer problems from the 2024 American Invitational Mathematics Examination I & II, directly sourced from the MAA’s public releases.
A.3.2 Evaluation Metrics
For the safety benchmarks, Sorry-bench, StrongREJECT, and WildJailBreak use attack successful rate (ASR) as the evaluation metric, revealing the times that a model accepts harmful prompts. Following the setting of the original benchmark, we used the rejection rate for JailBreakbench, measuring how often the model successfully rejects harmful prompts. For reasoning benchmarks, we use accuracy as the evaluation metric, measuring the rate at which models give correct answers.
A.3.3 Benchmark Hyperparameters Details
We used a rollout of 2 for GPQA-Diamond and 8 for AIME2024. The evaluations on GPQA-Diamond and AIME2024 were conducted using the OpenCompass framework. The detailed hyperparameter setting is shown in Table 12.
Appendix B Initial Investigation by Switching Thinking Mode
Recent works (Zhou et al., 2025a; Zhang et al., 2025b) have observed that LRMs tend to answer harmful questions. To further examine whether the thinking trajectory causally influences model safety, we evaluate multiple LRMs with the thinking mode switched on and off. Specifically, we test 2k harmful queries from WildJailbreak (Jiang et al., 2024a), and use Llama-Guard-3-8B (Llama Team, 2024) as the automatic safety judge.
Figure 7 reports the proportion of harmful answers under the two settings across Qwen3 variants. A consistent pattern emerges across model scales: enabling the thinking mode substantially increases the likelihood of producing harmful answers. For all three Qwen3 variants, turning on thinking leads to a higher harmful-answer rate, with increases from 35.4% to 41.4% (8B), 32.3% to 35.6% (14B), and 30.9% to 41.5% (32B).
Notably, this effect does not diminish with model scaling. In fact, the largest model exhibits the most pronounced gap, suggesting that stronger reasoning capacity alone does not guarantee safer behavior. Instead, explicit reasoning may create additional opportunities for models to rationalize or justify unsafe responses. Collectively, these results offer preliminary evidence that explicit reasoning trajectories may amplify safety risks, motivating a finer-grained investigation of how unsafe behavior is instantiated within intermediate reasoning steps rather than being determined solely by the final output.
Figure 7: Harmful answer rate (%) of Qwen3 models on 2k harmful queries from WildJailbreak, comparing thinking mode enabled vs. disabled. Safety is automatically judged by Llama-Guard-3-8B. Enabling thinking consistently increases the proportion of harmful answers across model scales.
Appendix C Additional Experimental Results
C.1 Analysis: PCA Analysis of 8B and 14B Models
We conducted a comparative analysis of models with different parameter sizes and fine-tuning methods (SafR and SafB), aiming to evaluate their impact on safety and representational clustering. The observed differences are visualized in Figure 8 and 9, while quantitative results across all configurations are reported in Table 7.
Figure 8: PCA of the Qwen3-8B representation space.
Figure 9: PCA of the Qwen3-14B representation space.
Model Scale
The 32B model consistently outperforms the 8B model across safety and clustering metrics. It exhibits a higher Safety Distance in all settings (Base, SafR, SafB), indicating better separation from harmful content. Its Silhouette Score at the Base stage (0.140) also exceeds that of the 8B model (0.120), reflecting a more structured internal representation.
Fine-Tuning Methods
Both Safety Recomposition (SafR) and Safety Backtrack (SafB) substantially improve model safety, as reflected by the centroid shifts in Figure 8 and Figure 9 and by the corresponding increases in Safety Distance reported in Table C.2. SafB reliably pushes the harmful cluster farther from the decision boundary while keeping the reasoning cluster relatively stable, indicating a balanced improvement in safety with limited impact on reasoning.
C.2 Preservation of Reasoning Token Usage
Table C.2 reports the average token length on GPQA-Diamond and AIME across training methods. Overall, SafR and SafB produce token counts comparable to the Base model, suggesting that their safety gains are not achieved by shortening generations. For example, on Qwen3-32B, the Base model averages 5150.81/12634.11 tokens (GPQA/AIME), while SafR remains close at 4830.90/12548.96 and SafB at 6452.00/12318.87. In contrast, SafePath shows a pronounced reduction to 3925.96/10200.49, with similar downward shifts on Qwen3-8B and Qwen3-14B. These results are consistent with the view that our methods improve safety while largely preserving the model’s intrinsic reasoning behavior.
Method GPQA-Diamond AIME
[…]
Base 7553.84 14895.49
SafePath 4869.04 11816.52
STAR-1 3449.49 12337.94
SafeChain 5105.35 12533.35
SafR 5212.79 14472.53
SafB 6540.14 14881.68
[…]
Base 5585.28 14419.94
SafePath 4379.39 10939.46
STAR-1 2736.43 11011.73
SafeChain 4688.32 12372.85
SafR 4303.19 13549.96
SafB 5091.59 13010.36
[…]
Base 5150.81 12634.11
SafeChain 5588.61 11198.09
SafePath 3925.96 10200.49
STAR-1 3564.03 11862.56
SafB 6452.00 12318.87
SafR 4830.90 12548.96
Table 13: Average token length on GPQA-Diamond and AIME benchmarks for Qwen3 models across different training methods.
Method Harmful $\Delta$ Reasoning $\Delta$
[…]
Base 12.996 0.000 13.983 0.000
SafB 19.615 +6.619 12.800 –1.183
SafR 19.615 +6.619 12.800 –1.183
SafeChain 13.609 +0.613 5.324 –8.659
[…]
Base 15.540 0.000 9.054 0.000
SafB 19.142 +3.602 9.632 +0.578
SafR 18.952 +3.412 8.328 –0.726
SafeChain 12.590 –2.951 4.265 –4.789
[…]
Base 11.197 0.000 12.180 0.000
SafB 15.158 +3.961 10.409 –1.771
SafR 16.182 +4.985 11.504 –0.676
SafeChain 11.576 +0.379 6.921 –5.259
Table 14: Safety distance and relative changes ($\Delta$) for Qwen3 models under different safety training methods.
Model S-B S-RB W-JB JB-B GPQA AIME
Qwen3-8B 45.45 13.62 38.80 18.29 57.58 73.30
R2D 19.32 0.96 12.80 84.15 41.92 47.92
SafR(Ours) 13.41 2.86 9.20 71.95 56.82 76.25
SafB(Ours) 17.05 2.05 8.00 73.17 54.30 77.50
Table 15: Performance comparison on safety and reasoning benchmarks between our COG and the R2D baseline. S-B denotes Sorry-bench, S-RB denotes StrongREJECT, W-JB denotes WildJailbreak, and JB-B denotes JailBreak Bench. Lower is better for S-B, S-RB, W-JB, and JB-B; higher is better for GPQA and AIME.
C.3 Preservation of Reasoning Token Usage
To further contextualize the safety–reasoning trade-off, we contrast two representative paradigms: (i) reasoning-for-safety (or reasoning-as-defense), which improves robustness by explicitly injecting safety-aware reasoning signals (e.g., reflection or step-level self-assessment), and (ii) safety-for-reasoning, which aims to enhance safety while preserving the integrity of an LRM’s original reasoning process.
Table 15 compares our COG-based variants with a representative reasoning-as-defense baseline, R2D. While R2D achieves strong gains on safety-related metrics (SafB/SafR and WildJailbreak-style benchmarks), it exhibits a substantial drop in reasoning performance on GPQA and AIME, indicating that heavy safety-oriented reasoning control may interfere with general reasoning quality. In contrast, our SafB/SafR maintain markedly stronger reasoning performance (GPQA/AIME) while remaining competitive on safety benchmarks, suggesting that trajectory-level interventions can better preserve reasoning behavior under safety constraints. Nevertheless, the results also indicate that fully preserving reasoning integrity while maximizing safety remains challenging, motivating further study on how safety constraints reshape intermediate reasoning trajectories.
Appendix D Case Study: COT Structure Stability
To further demonstrate that our method maintains structural consistency with the original model’s reasoning patterns, we conduct a case study on a problem from AIME 2024. As shown in Figure 10, SafR exhibits a highly consistent reasoning structure with Qwen3-8B, while Star-1 employs a fundamentally different approach and ultimately produces an incorrect answer.
Specifically, as highlighted in blue, both Qwen3-8B and SafR decompose the problem into two symmetric cases based on the starting direction, apply the same combinatorial formula $C(7,2)\times C(7,1)=147$ for each case, and correctly aggregate the results as $147+147=294$. In contrast, Star-1 (red highlights) introduces a "four types" classification strategy, leading to systematic double-counting and yielding $588=2\times 294$. This case illustrates that SafR preserves not only the base model’s reasoning steps but also its mathematical correctness.
Appendix E Example of Safety Failure in LRMs
E.1 Harm Misidentification
The example of Warning in Harm Misidentification is presented in Figure 14, where LRM answers the harmful quesiton as the regular one.
E.2 Self-Jailbreak: Benign Reframing
The example of Benign Reframing in Self-Jailbreak is presented in Figure 11, where red text represents their Self-Jailbreak act.
E.3 Self-Jailbreak: Logical Fallacies
The example of Logical Fallacies in Self-Jailbreak is presented in Figure 12, where red text represents their Self-Jailbreak act.
E.4 Self-Jailbreak: Warning
The example of Warning in Self-Jailbreak is presented in Figure 13, where red text represents their Self-Jailbreak act.
Appendix F Prompt Design
In this section, we present the prompt designs used in our COG framework and other experiments.
F.1 Extraction Prompt
This subsection provides both the base prompts and few-shot examples used in the extraction task.
Prompt
The basic prompt template used for extraction is shown in Figure 15.
Few-Shot
We also include few-shot examples to guide the model during extraction, as illustrated in Figure 16.
F.2 Classification Prompt
Prompt
The prompt used for the classification task is outlined below in Figure 17.
F.3 Safety Recomposition Prompt
For the Safety Recomposition stage, our prompt is constructed by concatenating several components. Specifically, based on the classification result, we extract a corresponding rewrite instruction from the “sub_prompts" field in a JSON file. Then, we combine the "main_prompt", the selected "sub_prompt", and a "format_prompt" to form the final prompt. The detailed structure is as shown in Figure 18
F.4 Safety Backtrack Prompt
In the Safety Backtrack stage, we further extend the prompt structure from the Safety Recomposition process. In addition to the previous components, we incorporate a transition phrase from “contextual_transition_phrases", selected based on the classification result. This helps guide the model more smoothly and maintain coherence in the final output. The detailed structure is as shown in Figure 19
F.5 Integration Prompt
This stage integrates the outputs from previous modules into a unified final response. The detailed prompt is as shown in Figure 20
Figure 10: Reasoning pattern comparison on an AIME 2024 problem (ground truth: 294). Blue highlights indicate where SafR and Qwen3-8B share identical reasoning structures. Red highlights show Star-1’s divergent approach leading to an incorrect answer (588).
Figure 11: Example of Benign Reframing.
Figure 12: Example of Warning.
Figure 13: Example of Warning.
Figure 14: Example of Harm Identification.
Figure 15: Example of the basic extraction prompt used in the extraction stage.
Figure 16: Few-shot examples used in the extraction prompt during the extraction stage.
Figure 17: Example of the basic classification prompt used in the classification stage.
Figure 18: Prompt structure used in the Safety Recomposition stage, formed by concatenating the main prompt, selected sub-prompt based on classification results, and format prompt.
Figure 19: Prompt structure used in the Safety Backtrack stage, incorporating contextual transition phrases based on classification results.
Figure 20: Prompt structure used in the Safety Backtrack stage, incorporating contextual transition phrases based on classification results.